OpenStudy (anonymous):

5. Evaluate the function f (x)  4 • 7x for x  1 and x = 2. Show your work.

OpenStudy (anonymous):

f(x) = 4 * 7x f(x) = 4 * 7(-1) Simplify that, and simplify: f(x) = 4 * (2)

OpenStudy (conqueror):

^^^^

OpenStudy (anonymous):

i still dont understand

OpenStudy (anonymous):

4x7=-28 4x2=8

OpenStudy (anonymous):

what are the boxes?

OpenStudy (anonymous):

evaluate the function f(x)=4x7^x for x=-1 and x=2. show your work @angiemarieeee

OpenStudy (anonymous):

x=2 would 196 i think..

OpenStudy (anonymous):

that is if you meant to put 7^x if it is just 7x then it would be 56

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

it is 7^x

OpenStudy (anonymous):

i have to solve it

OpenStudy (anonymous):

like evaluate the different quotient?

OpenStudy (anonymous):

i still dont know cause i have been in school since 2000 When I was in school I didnt understand it

OpenStudy (anonymous):

how though

OpenStudy (dan815):

okay what do u need help with

OpenStudy (anonymous):

0 evaluate the function f(x)=4x7^x for x=-1 and x=2. show your work

OpenStudy (dan815):

is it not \[4*7^x\]

OpenStudy (dan815):

\[f(x)=4*7^x\] when x is substituted with 2 you get \[f(2)=4*7^2=4*49=196\]

OpenStudy (dan815):

\[f(-1)=4*7^{-1}=4*1/7=4/7\]

OpenStudy (dan815):

when you have a negative exponent its the same as 1/that number to a positive exponent

OpenStudy (dan815):

for example

OpenStudy (dan815):

\[7^{-1} = \frac{1}{7^1}\] \[7^{-2} = \frac{1}{7^2}\] \[7^{-x} = \frac{1}{7^x}\]

OpenStudy (anonymous):

so is that the answers or an example

OpenStudy (anonymous):

@dan815 i still dont get it